OpenStudy (cutiecomittee123):

Find the exact value of tan 5pi/12

OpenStudy (lolacole12):

tan( 5π/12) = tan(3π/12 + 2π/12) = tan( π/4 + π/6) = [ tan(π/4) + tan(π/6) ] / [ 1 - tan(π/4)•tan(π/6)] = [ 1 + 1/√3 ] / [ 1 - 1 • 1/√3] = [ (3 + √3) /3 ] / [ (3 - √3) / 3] = (3 + √3) / (3 - √3) = (3 + √3)² / (9 - 3) = (9 + 6√3 + 3) / 6 = (12 + 6√3) / 6 = 2 + √3

OpenStudy (lolacole12):

medal